# JetFormer: an autoregressive generative model of raw images and text

### v1

@@ 2 Related Work
More recently, compressing the high-dimensional image pixel space to a lower-dimensional sequence of discrete tokens via a pretrained, frozen VQ-VAE (van den Oord et al., 2017; Razavi et al., 2019), and then modeling the compressed sequence with a transformer decoder has emerged as a scalable technique for high-fidelity image generation (Esser et al., 2020; Yu et al., 2022a; Ramesh et al., 2021; Yu et al., 2022c; Gafni et al., 2022). To enable semantic compression, VQ-VAEs typically rely on perceptual and GAN losses. Moreover, VQ-VAE-based representations are common in the context of dense prediction tasks (Kolesnikov et al., 2022; Lu et al., 2022; Mizrahi et al., 2024; Mentzer et al., 2024), in particular when modeling multiple modalities jointly. GIVT (Tschannen et al., 2024) showed that combining an autoencoder and an autoregressive transformer can be applied to continuous-valued sequences, by directly modeling feature vectors in the latent space of a VAE, without any quantization. Somewhat related, (Nachmani et al., 2023; Meng et al., 2024) explored soft tokens for speech synthesis.
@@ 4 Experiments
We rely on the simple, transformer-based design from (Anonymous, 2024) for the normalizing flow model. This design uses affine coupling layers (predicting an element-wise scale and shift) consisting of spatial and channel-wise splitting functions to split the activations in two parts, and a stack ViT blocks (Dosovitskiy et al., 2021) applied to half of the activations to infer the affine transform. Here, we only use channel-wise splitting as we found in preliminary experiments that spatial splitting did not improve modeling quality. We set the depth to 32 coupling blocks, each consisting of a stack of 4 or 6 ViT blocks of width 512, and with 8 attention heads. The (input and output) feature shape for this model is when feeding the full image, and when using a dense invertible map or PCA to reduce dimensionality prior to applying flow. For image size and patch size this amounts to , after flattening spatial dimensions and with dimensionality reduction to to .
@@ 4.1 Class-conditional image generation
Omitting the noise curriculum also results in much worse results.
@@
First training the normalizing flow to minimize NLL w.r.t. a Gaussian prior, and then training the autoregressive transformer on the latent representation of the frozen flow model leads to lower sample quality than end-to-end training.
Factoring out dimensions with learnable linear mapping before the flow leads to better results, but underperforms post-flow factoring out.
@@
Interestingly, we observe that modeling images after the PCA transform leads to somewhat worse results. However, in the presence of PCA, the noise curriculum becomes less important. It highlights the importance of prioritizing the important high-level information in images, if visual quality is the end goal. It also shows that manual preprocessing may reduce the necessity for various modeling tricks, yet it also shows that full end-to-end modeling prevails when done right.
@@ Appendix A Architectures and Hyperparameters
To implement the normalizing flow we use a stack affine coupling blocks (Dinh et al., 2016, Eqn. 7, 8). For each block the relation between input sequence and output is

### v2

@@ 2 Related Work
More recently, compressing the high-dimensional image pixel space to a lower-dimensional sequence of discrete tokens via a pretrained, frozen VQ-VAE (van den Oord et al., 2017; Razavi et al., 2019), and then modeling the compressed sequence with a transformer decoder has emerged as a scalable technique for high-fidelity image generation (Esser et al., 2020; Yu et al., 2022a; Ramesh et al., 2021; Yu et al., 2022c; Gafni et al., 2022). To enable semantic compression, VQ-VAEs typically rely on perceptual and GAN losses. Moreover, VQ-VAE-based representations are common in the context of dense prediction tasks (Kolesnikov et al., 2022; Lu et al., 2022; Mizrahi et al., 2024; Mentzer et al., 2024), in particular when modeling multiple modalities jointly. GIVT (Tschannen et al., 2024) showed that combining an autoencoder and an autoregressive transformer can be applied to continuous-valued sequences, by directly modeling feature vectors in the latent space of a VAE, without any quantization. Somewhat related, (Weiss et al., 2021; Nachmani et al., 2023; Meng et al., 2024) explored soft tokens for speech synthesis.
@@ 4 Experiments
We rely on the simple, transformer-based design from (Kolesnikov et al., 2024) for the normalizing flow model. This design uses affine coupling layers (predicting an element-wise scale and shift) consisting of spatial and channel-wise splitting functions to split the activations in two parts, and a stack ViT blocks (Dosovitskiy et al., 2021) applied to half of the activations to infer the affine transform. Here, we only use channel-wise splitting as we found in preliminary experiments that spatial splitting did not improve modeling quality. We set the depth to 32 coupling blocks, each consisting of a stack of 4 or 6 ViT blocks of width 512, and with 8 attention heads. The (input and output) feature shape for this model is when feeding the full image, and when using a dense invertible map or PCA to reduce dimensionality prior to applying flow. For image size and patch size this amounts to , after flattening spatial dimensions and with dimensionality reduction to to .
@@ 4.1 Class-conditional image generation
Omitting the noise curriculum further results in much worse results.
@@
First training the normalizing flow to minimize NLL w.r.t. a Gaussian prior, and then training the autoregressive transformer on the latent representation of the frozen flow model (factoring out redundant dimensions of the frozen latent space with a learnable, invertible linear map) results in lower sample quality than end-to-end training.
Factoring out dimensions with a learnable linear mapping before the flow (as described in Sec. 3.2) leads to better results, but underperforms post-flow factoring out.
@@
Interestingly, we observe that modeling images after the PCA transform leads to somewhat worse results. However, in the presence of PCA, the noise curriculum becomes less important. It highlights the importance of prioritizing the high-level information in images, if visual quality is the end goal. It also shows that manual preprocessing may reduce the necessity for various modeling tricks, yet it also shows that full end-to-end modeling prevails when done right.
@@ Appendix A Architectures and Hyperparameters
We rely on the design from (Kolesnikov et al., 2024) for the normalizing flow, which uses a stack affine coupling blocks (Dinh et al., 2016, Eqn. 7, 8). For each block the relation between input sequence and output is